QuestionCoveBot:

Calculate the slope between the two points. (5,2) and (0,−2) A) -5/4 B)-4/5 C) 4/5 D) 5/4

QuestionCoveBot:

@snowflake0531

snowflake0531:

Solve slope \[\frac{ y _{2} - y _{1} }{ x _{2} - x_{1} }\]

snowflake0531:

(x1,y1) (x2,y2)

QuestionCoveBot:

So it'd look like this? \[\frac{ -2-2 }{ 0-5 }\]

snowflake0531:

Yes, although it would be easier if you just switched it lol, then there would be less negative signs o-O

QuestionCoveBot:

Y is -4? X is 5?

snowflake0531:

like i mean \[\frac{ 2-(-2) }{ 5-0 }\]

snowflake0531:

@questioncovebot wrote:
Y is -4? X is 5?
wat

QuestionCoveBot:

-2-2 is -4 0-5 is 5 That's how I got -4 and 5

snowflake0531:

.... that whole thing is slope

snowflake0531:

@questioncovebot wrote:
-2-2 is -4 0-5 is 5 That's how I got -4 and 5
Also no the bottom would be -5, so there would be negative signs anymore just 4/5

QuestionCoveBot:

Oh so C ._. Thx snowflake

snowflake0531:

yw~